OpenStudy (anonymous):

Rationlize denominator and simplify. I get 102 with 3√6 + 5√2 / 4√6-3√2 Is this right?

OpenStudy (sepeario):

\[\frac{ 3\sqrt{6}+5\sqrt{2} }{ 4\sqrt{6} -3\sqrt{2}} * \frac{ { 4\sqrt{6} +3\sqrt{2}} }{ { 4\sqrt{6} +3\sqrt{2}} }\]

OpenStudy (sepeario):

I got 51+2 root 3.

OpenStudy (sepeario):

Fascinatingly half your answer.. :(

OpenStudy (anonymous):

Oh, yikes

OpenStudy (anonymous):

Would you be willing to stay and help me with this?

OpenStudy (anonymous):

I have other questions pertaining to this

OpenStudy (sepeario):

Give me two seconds please. I will work it out on paper and show you.

OpenStudy (anonymous):

Ok

OpenStudy (sepeario):

quickly checking did you multiply it with the conjugate surd like i did above?

OpenStudy (anonymous):

i dont remember lol

OpenStudy (anonymous):

still there?

OpenStudy (sepeario):

yeah

OpenStudy (anonymous):

ok. got the work? :)

OpenStudy (anonymous):

\[(51+29\sqrt3)\div39\] i think this should be you answer